If the local time of A is 7:30 am and longitude is 60 degrees west, what will the local time of town B be if it is 40 degrees east?
1
Expert's answer
2020-11-09T11:00:32-0500

The simple rule is that 15° correspond 1 hour difference. To the west of a prime meridian, time zones decrease, to the east time increases.

So, we have 60°+40°=100° degrees difference. In hours it will be


"T=\frac{100°}{15°/\text{h}}=\frac{20}{3}\text{ h},\text{ or }6\text{ h } 40\text{ min}."

Now, to find time at 40° E, all we need to do is to add 6 h 40 min to 7:30, which results in


"\tau=t+T=\bigg(7+\frac{30}{60}\bigg)+\bigg(6+\frac{40}{60}\bigg)=\frac{85}{6}\text{ h},\\\space\\
\tau=\frac{85}{6}=14+\frac16=14\text{ h } 10\text{ min}."

Thus, for 40°E, the time will be 2:10 pm.